### v1

@@
Abstract Sequence alignment is a cornerstone technique in computational biology for assessing similarities and differences among biological sequences. A key variant, sequence-to-graph alignment, plays a crucial role in effectively capturing genetic variations. In this work, we introduce two novel formulations within this framework: the Gap-sensitive Co-Linear Chaining (Gap-CLC) problem and the Co-Linear Chaining with Errors based on Edit Distance (Edit-CLC) problem, and we investigate their computational complexity. We show that solving the Gap-CLC problem in sub-quadratic time is highly unlikely unless the Strong Exponential Time Hypothesis (SETH) fails—even when restricted to binary alphabets. Furthermore, we establish that the Edit-CLC problem is NP-hard in the presence of errors within the graph. These findings emphasize that incorporating ”co-linear” structures into sequence-to-graph alignment models fails to reduce computational complexity, highlighting that these models remain at least as computationally challenging to solve as those lacking such prior information.
Keywords Co-linear chaining; Sequence alignment; Algorithm; Complexity
@@ 1. Introduction
Given the vast amount of sequence data and the need to evaluate numerous possible alignments, sequence alignment is inherently computationally demanding. To address this challenge, researchers have introduced efficient optimization techniques and heuristic strategies. These include dynamic programming [29], which guarantees optimal solutions by breaking the problem into smaller subproblems, the widely used BLAST (Basic Local Alignment Search Tool) [16] that accelerates searches through heuristic filtering, and the sensitive Smith-Waterman algorithm [28] for precise local alignment. These methods collectively enhance both the accuracy and efficiency of sequence analysis.
Sequence-to-sequence alignment, naturally extending from classical sequence alignment, has been extensively studied since the 1970s. Dynamic programming methods [29] were among the first to address this problem. Myers and Miller [21] proposed a multiple-sequence alignment algorithm based on the line-sweep paradigm, leveraging range trees for orthogonal range searching to identify the highest-scoring alignment, though with a time complexity increased by a logarithmic factor. In 2005, Abouelhoda and Ohlebusch [1] introduced improved line-sweep algorithms that efficiently solve both global and local fragment-chaining problems in multiple genomes, reducing both time and space complexity compared to earlier approaches. Recent years have seen enhanced models incorporating gaps and overlaps in sequence-to-sequence alignment [4, 18, 12].
@@
In recent years, both exact and approximation algorithms for sequence-to-graph alignment have received significant attention. When the alignment path is constrained to visit distinct vertices, the problem becomes NP-hard [17] on general pangenome graphs, including De Bruijn graphs [11], as it encompasses the classical Hamiltonian Path problem. However, most existing approaches relax this constraint. Sequence-to-graph alignment on general graphs was initially explored two decades ago in the context of hypertext search [3], where Amir and Lewenstein proposed an $O(|V|+m|E|)$-time algorithm for exact alignment, with $m$ denoting the pattern length. For restricted graph classes, efficient solutions exist: polynomial-time algorithms apply to DAGs [24] and trees [2], with time complexities of $O(|V|+m|E|)$ and $O(|V|)$, respectively. Recently, Equi et al. [9] established a conditional lower bound, demonstrating that, for any constant $\gamma>0$, unless the Strong Exponential Time Hypothesis (SETH) fails, no algorithm can solve the exact alignment problem in $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time on general graphs—even over a binary alphabet.
@@
Nevertheless, co-linear chaining in sequence-to-graph alignment remains more complex than in sequence-to-sequence alignment, posing greater computational challenges. Mäkinen et al. [19] investigated co-linear chaining on directed acyclic graphs (DAGs), developing a sparse dynamic programming approach tailored to DAGs. Subsequently, gap costs were incorporated into the scoring function [6]. More recently, Rajput et al. [25] proposed a practical formulation and an efficient algorithm for co-linear chaining on general graphs, extending gap cost considerations to both query sequences and pan-genome graphs.
@@
  complexity  analyzes the computational complexity of both problems. Intuitively, models ignoring ”co-linear” structures can be reduced to the Gap-CLC and Edit-CLC frameworks, implying that incorporating co-linearity does not reduce complexity and may in fact preserve or increase it. We conclude with a summary and discussion of future directions in Section
  conclude  .
@@ 2.1 Problem formulations
Problem 1: Gap-sensitive Co-linear Chaining problem (Gap-CLC problem)
@@
Problem 2: Co-linear Chaining with errors based on Edit distance (Edit-CLC problem)
@@ 3. The complexity
In this section, each vertex in the graph is assigned a single-character label, which is sufficient to demonstrate the inherent difficulty of the Gap-CLC and Edit-CLC problems. We present a linear-time reduction from the Single-Exa-SGM problem to Gap-CLC, establishing that no sub-quadratic time algorithm exists for Gap-CLC, unless the Single-Exa-SGM problem admits such an algorithm, which is ruled out even over a binary alphabet [9]. Specifically, for any constant $\gamma>0$, the Single-Exa-SGM problem cannot be solved in either $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time unless the Strong Exponential Time Hypothesis (SETH) fails, where $E$ denotes the edge set of the graph and $m$ is the length of the query sequence [9].
@@
There is no sub-quadratic time algorithm for the Gap-CLC problem, unless the Strong Exponential Time Hypothesis (SETH) is false, even the alphabet is binary.
@@
Consequently, we have developed a sub-quadratic time algorithm for the Single-Exa-SGM problem, challenging the recent assertion that no such algorithm exists unless the Strong Exponential Time Hypothesis (SETH) fails—even over a binary alphabet [9]. This implies that the Gap-CLC problem also admits no sub-quadratic time algorithm.
@@
4. Conclusions and discussions
Co-linear chaining problems have found extensive applications in sequence-to-sequence alignment, yet their extension to sequence-to-graph alignment presents greater challenges. In this work, we investigate the computational complexity of co-linear chaining in the context of sequence-to-graph alignment, focusing specifically on the Gap-CLC and Edit-CLC problems. We demonstrate that the Gap-CLC problem is at least as hard as the Exa-SGM problem, while the Edit-CLC problem is proven to be NP-hard when errors are permitted on pan-genome graphs—even over a binary alphabet. These complexity results reveal that incorporating the ”co-linear” constraint in sequence-to-graph alignment does not reduce the theoretical complexity compared to non-co-linear models. Despite these theoretical challenges, ”co-linear” chaining remains widely adopted in practical tools, highlighting a significant gap between theoretical understanding and real-world practice. Bridging this gap presents an important and intriguing direction for future research. Furthermore, algorithmic solutions for the Edit-CLC problem remain scarce, with limited progress on both exact and approximation algorithms, indicating a clear need for further investigation and development in this area.

### v2

@@
Abstract Sequence alignment is a cornerstone technique in computational biology for assessing similarities and differences among biological sequences. A key variant, sequence-to-graph alignment, plays a crucial role in effectively capturing genetic variations. In this work, we introduce two novel formulations within this framework: the Gap-sensitive Co-Linear Chaining (Gap-CLC) problem and the Co-Linear Chaining with Errors based on Edit Distance (Edit-CLC) problem, and we investigate their computational complexity. We show that solving the Gap-CLC problem in sub-quadratic time is highly unlikely unless the Strong Exponential Time Hypothesis fails—even when restricted to binary alphabets. Furthermore, we establish that the Edit-CLC problem is NP-hard in the presence of errors within the pan-genome graph. These findings emphasize that incorporating co-linear structures into sequence-to-graph alignment models fails to reduce computational complexity, highlighting that these models remain at least as computationally challenging to solve as those lacking such prior information.
Keywords Sequence-to-graph alignment; Co-linear chaining; Gap-sensitive Co-Linear Chaining; Edit Distance Co-Linear Chaining; Computational complexity; Strong Exponential Time Hypothesis; NP-hard
@@ 1. Introduction
Given the vast amount of sequence data and the need to evaluate numerous possible alignments, sequence alignment is inherently computationally demanding. To address this challenge, researchers have introduced efficient optimization techniques and heuristic strategies. These include dynamic programming [25], which guarantees optimal solutions by breaking the problem into smaller subproblems, the widely used Basic Local Alignment Search Tool [16] that accelerates searches through heuristic filtering, and the sensitive Smith-Waterman algorithm [29] for precise local alignment. These methods collectively enhance both the accuracy and efficiency of sequence analysis.
Sequence-to-sequence alignment, naturally extending from classical sequence alignment, has been extensively studied since the 1970s. Dynamic programming methods [25] were among the earliest to address this problem. Myers and Miller [21] proposed a multiple-sequence alignment algorithm based on the line-sweep paradigm, leveraging range trees for orthogonal range searching to identify the highest-scoring alignment, though with a time complexity increased by a logarithmic factor. In 2005, Abouelhoda and Ohlebusch [1] introduced improved line-sweep algorithms that efficiently solve both global and local fragment-chaining problems in multiple genomes, reducing both time and space complexity compared to earlier approaches. Recent years have seen enhanced models incorporating gaps and overlaps in sequence-to-sequence alignment [4, 18, 12].
@@
In recent years, both exact and approximation algorithms for sequence-to-graph alignment have received significant attention. When the alignment path is constrained to visit distinct vertices, the problem becomes NP-hard [17] on general pangenome graphs, including De Bruijn graphs [11], as it encompasses the classical Hamiltonian Path problem. However, most existing approaches relax this constraint. Sequence-to-graph alignment on general graphs was initially explored two decades ago in the context of hypertext search [3], where Amir and Lewenstein proposed an $O(|V|+m|E|)$-time algorithm for exact alignment, with $m$ denoting the pattern length. For restricted graph classes, efficient solutions exist: polynomial-time algorithms apply to Directed Acyclic Graphs (DAGs) [24] and trees [2], with time complexities of $O(|V|+m|E|)$ and $O(|V|)$, respectively. Recently, Equi et al. [9] established a conditional lower bound, demonstrating that, for any constant $\gamma>0$, unless the Strong Exponential Time Hypothesis (SETH) fails, no algorithm can solve the exact alignment problem in $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time on general graphs—even over a binary alphabet.
@@
Nevertheless, co-linear chaining in sequence-to-graph alignment remains more complex than in sequence-to-sequence alignment, posing greater computational challenges. Mäkinen et al. [19] investigated co-linear chaining on DAGs, developing a sparse dynamic programming approach tailored to DAGs. Subsequently, gap costs were incorporated into the scoring function [6]. More recently, Rajput et al. [26] proposed a practical formulation and an efficient algorithm for co-linear chaining on general graphs, extending gap cost considerations to both query sequences and pan-genome graphs.
@@
  complexity  analyzes the computational complexity of both problems. Intuitively, models ignoring ”co-linear” structures can be reduced to the Gap-CLC and Edit-CLC frameworks, implying that incorporating co-linearity does not reduce complexity and may in fact preserve or increase it. We conclude with a summary in Section
  conclude  and propose some future work in Section
  discussions  .
@@ 2.1 Problem Formulations
Problem 1: Gap-CLC problem
@@
Problem 2: Edit-CLC problem
@@ 3. The Complexity
In this section, each vertex in the graph is assigned a single-character label, which is sufficient to demonstrate the inherent difficulty of the Gap-CLC and Edit-CLC problems. We present a linear-time reduction from the Single-Exa-SGM problem to Gap-CLC, establishing that no sub-quadratic time algorithm exists for Gap-CLC, unless the Single-Exa-SGM problem admits such an algorithm, which is ruled out even over a binary alphabet [9]. Specifically, for any constant $\gamma>0$, the Single-Exa-SGM problem cannot be solved in either $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time unless the SETH fails, where $E$ denotes the edge set of the graph and $m$ is the length of the query sequence [9].
@@
There is no sub-quadratic time algorithm for the Gap-CLC problem, unless the SETH is false, even the alphabet is binary.
@@
Consequently, we have developed a sub-quadratic time algorithm for the Single-Exa-SGM problem, challenging the recent assertion that no such algorithm exists unless the SETH fails—even over a binary alphabet [9]. This implies that the Gap-CLC problem also admits no sub-quadratic time algorithm.
@@
4. Conclusions
Co-linear chaining problems have found extensive applications in sequence-to-sequence alignment, yet their extension to sequence-to-graph alignment presents greater challenges. In this work, we investigate the computational complexity of CLC in the context of sequence-to-graph alignment, focusing specifically on the Gap-CLC and Edit-CLC problems. We demonstrate that the Gap-CLC problem is at least as hard as the Exa-SGM problem, while the Edit-CLC problem is proven to be NP-hard when errors are permitted on pan-genome graphs—even over a binary alphabet. These complexity results reveal that incorporating the ”co-linear” constraint in sequence-to-graph alignment does not reduce the theoretical complexity compared to non-co-linear models.
5. Discussions
Despite these theoretical challenges, ”co-linear” chaining remains widely adopted in practical tools, highlighting a significant gap between theoretical understanding and real-world practice. Bridging this gap presents an important and intriguing direction for future research. Furthermore, algorithmic solutions for the Edit-CLC problem remain scarce, with limited progress on both exact and approximation algorithms, indicating a clear need for further investigation and development in this area.